OpenStudy (mackenzie2013):

Fifty-four is __% of 60.

OpenStudy (mathstudent55):

Write an equation and solve it. Fifty-four is __% of 60. 54 = x% * 60 % means divide by 100 54 = x/100 * 60 54 = 60x/100 Now solve for x.

OpenStudy (anonymous):

it's 90

OpenStudy (mackenzie2013):

So you would divide 60 by 100?

OpenStudy (anonymous):

no i would multiply by 100 both side...

OpenStudy (jdoe0001):

if we take 60 as being the 100% then one could say that \(\large \begin{array}{llll} quantity&percetage\\ \hline\\ 60&100\\ x&54 \end{array}\implies \bf \cfrac{60}{x}=\cfrac{100}{54}\)

OpenStudy (jdoe0001):

darn typos... but anyow, =)

OpenStudy (jdoe0001):

\(\large \begin{array}{ccllll} quantity&percentage\\ \hline\\ 60&100\\ x&54 \end{array}\implies \bf \cfrac{60}{x}=\cfrac{100}{54}\) so you'd solve for "x"

OpenStudy (jdoe0001):

hmmmm shoot... ...lemme fix that... 54 is a quantity

OpenStudy (jdoe0001):

\(\large \begin{array}{ccllll} quantity&percentage\\ \hline\\ 60&100\\ 54&x \end{array}\implies \bf \cfrac{60}{54}=\cfrac{100}{x}\) there

OpenStudy (mackenzie2013):

So you would cross multiply, correct?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

this one man